OpenStudy (anonymous):

In 2003, the population of Charlotte, NC was 584,658. According to the U.S. Census Bureau, Charlotte has been growing at the rate of 4.8% annually since 2003. What equation models the population of Charlotte t years after 2003?

OpenStudy (anonymous):

And explain why please. A y = 584,658(1 + 4.8)t B y = 584,658(1 + 0.48)t C y = 584,658(1 + 0.048)t D y = 584,658(0.048)t

OpenStudy (kropot72):

After one year the population will be given by: \[\large Pop.\ after\ 1\ year=584,658\times1.048\] In the choices, if you substitute 1 for t, which choice matches the above expression?

OpenStudy (wolf1728):

I am wondering if all those choices should end with ^t instead of *t?

OpenStudy (kropot72):

@wolf1728 Yes, you are correct. Good spotting!

OpenStudy (wolf1728):

Thanks kropot :-)

OpenStudy (kropot72):

yw

OpenStudy (kropot72):

If Pt is the population after t years, P0 is the initial population and r is the growth rate as a decimal the general equation is: \[\large P _{t}=P _{0}(1+r)^{t}\] Hopefully you can make the correct choice now.